The given system of linear equations has a unique solution
2x + y- z = -5
-5x - 3y + 2z = 7
x + 4y - 3z = 0
what is the value of y?

Answer by rothauserc(4718)

Determinant = 6
:
Dx = -12, Dy = 30, Dz = 36
:
x = -12/6 = -2
y = 30/6 = 5
z = 36/6 = 6
:
**********************
x = -2, y = 5, z = 6
**********************
: